# WildFrame: Comparing Framing in Humans and LLMs on Naturally Occurring Texts

### v1

@@ 5 Limitations
Additionally, our approach to framing is highly specific. We only manipulate statements by adding a prefix or suffix, whereas reframing can take many other forms, such as restructuring sentences or altering word choices to convey ambiguous sentiment. This may limit the generalizability of our results.
@@ Appendix B Adding Framing
Base Sentence Base Sentiment Opposite Framing Sentence
“To my surprise I did enjoy the book and the characters .” Positive
“To my surprise I did enjoy the book and the characters, even though
it had a disappointing ending. ”
“For me , Aspartame causes bad memory loss and nasty
gastrointestinal distress .”
Negative
“For me, Aspartame causes bad memory loss and nasty gastrointestinal
distress, but this has encouraged me to seek out healthier, natural
alternatives and cultivate a balanced diet .”
Table 2: Sentences after framing. Positive sentences are added with negative framing, and vice-versa. The opposite framing is in bold.
Example for statements after framing are presented in in Table 2.
B.1 Framing Prompts
  1. 1.
    “Here is an example of a base statement with a negative sentiment: I failed my math test today. Here is the same statement, after adding a positive framing: I failed my math test today, however I see it as an opportunity to learn and improve in the future. Here is a negative statement: <statement> Like the example, add a positive suffix or prefix to it. Don’t change the original statement.”
  2. 2.
    “Here is an example of a base statement with a positive sentiment: I got an A on my math test. Here is the same statement, after adding a negative framing: I got an A on my math test. I think I spent too much time learning to it though. Here is a positive statement: <statement>. Like the example, add a negative suffix or prefix to it. Don’t change the original statement.”
@@ Appendix C Annotation Platform
Figure 5: Screenshot of the annotation platform.
@@ Appendix D Models
  • "meta-llama/Llama-3-8b-chat-hf"
  • "meta-llama/Llama-3-70b-chat-hf"
For GPT-4o, we used the OpenAI api, with "gpt-4o-2024-08-06". [^7] https://platform.openai.com/docs/overview
@@
Figure 7: Pairwise Pearson correlation coefficients between predictions from different models, indicating the degree of similarity in their behavior under opposite sentiment framing scenarios.

### v2

@@ Appendix B Adding Framing
We used three different methods to generate the reframings.
B.1 Reframing with GPT-4
These reframings were shown to human annotators and are the main reframings we analyze in this paper.
We used the gpt-4-0613 version for generating the reframing, using the following prompts:
  • “Here is an example of a base statement with a negative sentiment: I failed my math test today. Here is the same statement, after adding a positive framing: I failed my math test today, however I see it as an opportunity to learn and improve in the future. Here is a negative statement: STATEMENT Like the example, add a positive suffix or prefix to it. Don’t change the original statement.”
  • “Here is an example of a base statement with a positive sentiment: I got an A on my math test. Here is the same statement, after adding a negative framing: I got an A on my math test. I think I spent too much time learning it though. Here is a positive statement: STATEMENT. Like the example, add a negative suffix or prefix to it. Don’t change the original statement.”
B.2 Paraphrasing the reframings with Gemini-2.5-flash
Next, we utilized Gemini-2.5-flash Comanici et al. (2025) to paraphrase the already-reframed sentences, while preserving factual content but varying style, order, and wording.
We used the following prompt:
  • “Given this original sentence: STATEMENT, a former process has generated a reframing sentence. Your task is to rewrite the reframing, while maintaining its meaning and keep the same facts, but reframe it in a different way. Rewrite the following sentence using a different style or structure (e.g., change the tone, word order, or phrasing) while strictly preserving the original meaning and factual details.
    Original Sentence: STATEMENT.
    Original Reframing Sentence: REFRAMED.
    New Reframing:”
B.3 Paraphrasing given only the base statement
Finally, we want to generate a completly new reframing, that would introduce different information than in the first reframing. For that, we prompted Gemini-2.5-flash with only the base statement, and the cognitive definition of framing. We used the following prompt:
  • “In psychology and communication, Framing refers to the way information is presented to an audience. Specifically, Attribute Framing involves highlighting different characteristics of a single event to shift the perceived value from positive to negative (or vice versa), while the underlying facts of the event remain objectively true. Your task is to rewrite the provided BASE_SENTIMENT statement to apply OPPOSITE_SENTIMENT sentiment frame, i.e., FRAMING_PROMPT. You have full creative freedom to restructure the sentence, change the vocabulary, or alter the syntax, as long as you preserve the core factual event. Do not invent new facts, but you may interpret the existing facts through a new lens.
    Statement: STATEMENT
    Re-framed sentence:”
  • Positive FRAMING_PROMPT: “rewrite it to emphasize the positive implications (opportunity, safety, freedom, or long-term benefit)”
  • Negative FRAMING_PROMPT: “rewrite it to emphasize the negative implications (responsibility, risk, cost, or pressure).”
@@ Appendix C Annotation Platform
Figure 8: Screenshot of the Mechanical Turk annotation platform.
@@ Appendix D Models
  • "meta-llama/Llama-3.1-8B-Instruct-Turbo"
  • "meta-llama/Llama-3.3-70B-Instruct-Turbo"
For GPT model family, we used the OpenAI api, with “gpt-3.5-turbo-0125”, “gpt-4o-2024-08-06”, and “gpt-5-mini”. [^7] https://platform.openai.com/docs/overview
For Gemini-2.5-flash, we used Google AI Studio API, calling “gemini-2.5-flash”. [^8] https://aistudio.google.com
Appendix E In-House Annotations
We used a streamlit environment to run the in-house annotation exxperiment. We had 8 participants, all are grad students that agreed to participate in our experiment without payment.
In Figure 9 we provide a screenshot of the annotation platform, and their instructions for the task were similar to the Mechanical Turk experiment.
Each annotator has annotated random unique sentences from WildFrame, about 20 sentences each, resulting in a total of 160 in-house annotations which we know did not involve LLM usage.
Appendix F Dataset License and Usage Terms
The WildFrame dataset is released to support academic research on framing effects in humans and large language models. The dataset is distributed under a Creative Commons Attribution 4.0 International (CC BY 4.0) license.
Under this license, users are permitted to share, copy, redistribute, and adapt the dataset for any purpose, including commercial use, provided that appropriate credit is given to the original authors. Attribution should include a citation to this paper and a reference to the dataset name, WildFrame.
Source Data and Derivative Content
The WildFrame dataset is constructed from short excerpts of the Amazon Reviews dataset, which is publicly available and widely used for research purposes. Base statements were extracted using syntactic patterns and further filtered through human annotation. Reframed variants were generated using a large language model and therefore constitute derived textual content rather than verbatim reproductions of the original reviews.
The dataset does not include user identifiers, personal data, or metadata that could be used to identify individuals. All examples are presented as standalone sentences without links to original reviews or authors.
Intended Use and Limitations
The dataset is intended for research and evaluation purposes, including the study of framing effects, sentiment ambiguity, and human–model behavioral alignment. While the CC BY 4.0 license permits broad reuse, users are responsible for ensuring that their downstream applications comply with applicable laws, ethical guidelines, and the terms of any third-party models used in conjunction with the dataset.
We encourage responsible use of WildFrame, particularly in sensitive application domains, and recommend clearly documenting any modifications or extensions made to the dataset in derivative works.
Annotator ID Human % LLM % Diff (Human - LLM) # Annotations
u103c35 75.56% 76.67% -1.11% 90
u1806a5 71.64% 64.18% +7.46% 67
u1cafca 64.52% 58.06% +6.45% 31
u7b0b64 72.88% 67.80% +5.08% 59
u941c06 67.12% 68.49% -1.37% 73
u96efbf 60.00% 57.14% +2.86% 35
ua1ba86 79.85% 82.84% -2.99% 134
uc57e9e 81.10% 72.44% +8.66% 127
uf993a0 80.52% 80.52% +0.00% 154
In-House 66.23% 68.18% -1.95% 154
Table 3: Individual annotator agreement with human consensus (calculated via leave-one-out, including the in-house annotation) versus LLM Consensus. Negative values in the Diff column indicate higher agreement with the LLM cluster, while positive values indicate higher agreement with the human consensus. The annotation counts represent the subset of the 160 in-house control items labeled by each worker, excluding instances where the LLM ensemble produced a tie.
Figure 9: Screenshot of the in-house annotation platform via streamlit.
(a) Paraphrasing the reframing.
(b) Generating a new reframing from scratch.
Figure 10: Similar to Figure 4, percentage of reframed statements that results in sentiment shift positive to negative or vice versa), on two additional reframing paraphrases.
@@ Appendix F Dataset License and Usage Terms
(a) GPT-3.5
(b) GPT-4o
(c) GPT-5-mini
(d) Mistral-v0.3-7B
(e) Mixtral-v0.1-8x7B
(f) Mixtral-v0.1-8x22B
(g) Gemma-2-9B
(h) Gemma-2-27B
(i) Gemini-2.5-flash
(j) Llama-3.1-8B
(k) Llama-3.3-70B
Figure 12: Consistency under prompt paraphrasing per model, row normalized, corresponds to Figure 6.